OpenStudy (anonymous):

A force of 4 N acts on a body of mass 2 kg for 4 s. Assuming that the body to be initially at rest find : A) its velocity when force stops acting on it. B) The distance covered in 10 s after force starts acting on it

OpenStudy (anonymous):

Did you solve it?

OpenStudy (anonymous):

F*t =m(v-u) 4 * 4 = 2 (v-0) 16 = 2v v = 8 b) F = m x a 4 = 2a a=2 s= ut plus a/2at^2 s = 0 plus 1/2 x 2 x 10^2 s= 100